### v2

@@ 5.5 Parameter Identifiability through Stochasticity
is unimodal, exhibiting a clear and unique optimum. Otherwise, the parameter is interpreted as practically non-identifiable, in the sense that multiple values of the input may yield the same model output. We compute the posterior distribution using the Metropolis Hastings algorithm, a Markov chain Monte Carlo (MCMC) method using delayed rejection adaptive metropolis (DRAM) algorithm developed in [29]. The prior distribution $p^{prior}(\theta)$ of each parameter is taken to be a uniform distribution on the range in Table 3, assuming that we have no prior knowledge about the parameter other than its range. We choose a Gaussian distribution for the likelihood function $p(D|\theta)$ as following,
$\mathbf{p(D|\theta)=\prod_{i\in I}\frac{1}{\sqrt{2\pi\sigma^{2}}}{\exp}\left(\frac{(D_{i}-F(N_{i};\theta))^{2}}{2\sigma^{2}}\right)}$
where $F(N_{i};\theta)$ are the birth and death rate functions in Equations (2829, 3233) to be fitted that depend on parameters $\theta$. This form of likelihood assumes an additive error that is Gaussian. We run 8 independent runs of MCMC with a burn-in period of 10000 steps, followed by 50000 iterations. The initial points are chosen using Latin hypercube samples [30]. The convergence of the MCMC algorithm is checked via $\hat{R}$ [31] and the value is found to be 1.003 (significantly below the typical tolerance 1.05).
In Figure 10, prior distribution $p^{prior}(\theta)$ and posterior distributions $p^{post}(\theta|D)$ are plotted. The posterior distributions of all parameters are unimodal around the true parameter values, indicating accuracy of our inference method. In addition, Table 2 shows the statistics of the fitted parameter values for PC3 cell line. The median and 90% confidence interval drawn from the posterior distribution of each parameter are close to the true parameter value. This suggests that the proposed inference method is practically identifiable, and it can estimate the birth and death process parameters $\gamma_{k}$, $\sigma_{k}$, and $\delta_{k}$. We note that our inference method and results include a mix of frequentist and Bayesian methods. While the statistical properties of the method require further theoretical validation, simulation results demonstrate its practical effectiveness.
Figure 10: Prior (dash) and posterior (line)
probability distributions of the parameters, computed using a Markov chain Monte Carlo method (MCMC). The posterior distribution of each parameters is a unimodal function centered around the true parameter value (vertical dash-dot line). This suggests that our inference method is practically identifiable, and it is able to uniquely estimate the parameters that distinguish the effects of the birth and death processes.
@@
Table 2: Estimated parameters compared to the true parameter values. Considering the median and 90% confidence interval drawn from the posterior distribution, the inferred parameters are close to the true value in PC3 cell line [10]. See Appendix A for additional information.
@@ 6 Conclusions and Discussion
In the Lotka-Volterra model, we see several notable impacts on the way the subpopulations evolve when changing how the intraspecies and interspecies interactions affect the birth and death rates. Moving the intraspecies competition to the birth rate decreases the range of the population sizes seen after a long enough period of time. Changing the interspecies interaction from the birth rate to the death rate does not have as big of an impact on the final subpopulation sizes. This is true for both the PC3 and the DU145 cell line parameter values. The placement of the interspecies interaction in the birth versus the death rate has a much bigger impact on the survival probability of the resistant population. The type of interaction determines whether the survival probability increases or decreases as the interspecies interaction moves from the death rate to the birth rate. The intraspecies competition has less of an impact on the survival probability. This is likely because when the resistant subpopulation size is small and therefore has a reasonable chance of extinction, the size of the intraspecies term is dwarfed by the size of the interspecies interaction term. Once the resistant population has grown to a size where the intraspecies term is comparable to the interspecies term, the probability of the population going extinct is small enough to be negligible on the time scale of the calculations.
We introduced an inference method for disambiguating the birth and the death rate in a general birth-death process with $n$ distinguishable subtypes. The method requires several different time series of the counts of each subpopulation over time. This method introduced in Section 5.2 can be used even if the functional forms of the birth rate and the death rate are unknown. The main idea behind the method is to utilize the properties of the stochastic system to infer the rates. Here, we use the mean and variance, which is similar to the idea of the method of moments inference technique, which uses moments of the data to infer different parameter values. One could consider using a maximum likelihood approach to infer these parameters as well. However, because the dataset does not contain the birth and death counts, only the population counts at different time points, the distribution needed for a maximum likelihood analysis is that of the difference between two Poisson random variables, known as the Skellam distribution. An analysis comparing the method of moments to maximum likelihood techniques in this setting was done by Alzaid et al. 2010 [32].

### v3

@@ 5.5 Parameter Identifiability through Stochasticity
is unimodal, exhibiting a clear and unique optimum. Otherwise, the parameter is interpreted as practically non-identifiable, in the sense that multiple values of the input may yield the same model output. We compute the posterior distribution using the Metropolis Hastings algorithm, a Markov chain Monte Carlo (MCMC) method using the delayed rejection adaptive metropolis (DRAM) algorithm developed in [29]. The prior distribution $p^{prior}(\theta)$ of each parameter is taken to be a uniform distribution on the range in Table 3, assuming that we have no prior knowledge about the parameter other than its range. We choose a Gaussian distribution for the likelihood function $p(D|\theta)$,
$p(D|\theta)=\prod_{i\in I}\frac{1}{\sqrt{2\pi\sigma^{2}}}{\exp}\left(\frac{(D_{i}-F(N_{i};\theta))^{2}}{2\sigma^{2}}\right)$ (35)
where $F(N_{i};\theta)$ are the birth and death rate functions in Equations (2829, 3233) to be fitted, which depend on parameters $\theta$. This form of the likelihood assumes an additive error that is Gaussian. We run 8 independent runs of MCMC with a burn-in period of 10000 steps, followed by 50000 iterations. The initial points are chosen using Latin hypercube samples [30]. The convergence of the MCMC algorithm is checked via $\hat{R}$ [31] and the value is found to be 1.003 (significantly below the typical tolerance 1.05).
@@
Table 2: Estimated parameters compared to the true parameter values. Considering the median and 90% confidence interval drawn from the posterior distribution, the inferred parameters are close to the true value in the PC3 cell line [10]. See Appendix A for additional information.
In Figure 10, the prior distribution $p^{prior}(\theta)$ and posterior distributions $p^{post}(\theta|D)$ are plotted. The posterior distributions of all parameters are unimodal around the true parameter values, indicating accuracy of our inference method. In addition, Table 2 shows the statistics of the fitted parameter values for the PC3 cell line. The median and 90% confidence interval drawn from the posterior distribution of each parameter are close to the true parameter value. This suggests that the proposed inference method is practically identifiable, and it can estimate the birth and death process parameters $\gamma_{k}$, $\sigma_{k}$, and $\delta_{k}$. We note that our inference method and results include a mix of frequentist and Bayesian methods. While the statistical properties of the method require further theoretical validation, simulation results demonstrate its practical effectiveness.
Figure 10: Prior (dash) and posterior (line) probability distributions of the parameters, computed using a Markov chain Monte Carlo method (MCMC). The posterior distribution of each parameter is a unimodal function centered around the true parameter value (vertical dash-dot line). This suggests that our inference method is practically identifiable, and it is able to uniquely estimate the parameters that distinguish the effects of the birth and death processes.
@@ 6 Conclusions and Discussion
In this paper, we study the significance of ecological interactions and separation of birth and death dynamics in stochastic heterogeneous populations via general birth-death processes. In the Lotka-Volterra model, we see several notable impacts on the way the subpopulations evolve when changing how the intraspecies and interspecies interactions affect the birth and death rates. Moving the intraspecies competition to the birth rate decreases the range of the population sizes seen after a long enough period of time. Changing the interspecies interaction from the birth rate to the death rate does not have as big of an impact on the final subpopulation sizes. This is true for both the PC3 and the DU145 cell line parameter values. The placement of the interspecies interaction in the birth versus the death rate has a much bigger impact on the survival probability of the resistant population. The type of interaction determines whether the survival probability increases or decreases as the interspecies interaction moves from the death rate to the birth rate. The intraspecies competition has less of an impact on the survival probability. This is likely because when the resistant subpopulation size is small and therefore has a reasonable chance of extinction, the size of the intraspecies term is dwarfed by the size of the interspecies interaction term. Once the resistant population has grown to a size where the intraspecies term is comparable to the interspecies term, the probability of the population going extinct is small enough to be negligible on the time scale of the calculations.
We introduced an inference method for disambiguating the birth and the death rate in a general birth-death process with $n$ distinguishable subtypes. Our inference method requires several different time series of the counts of each subpopulation over time. This method introduced in Section 5.2 can be used even if the functional forms of the birth rate and the death rate are unknown. The main idea behind the method is to utilize the properties of the stochastic system to infer the rates. Here, we use the mean and variance, which is similar to the idea of the method of moments inference technique, which uses moments of the data to infer different parameter values. One could consider using a maximum likelihood approach to infer these parameters as well. However, because the dataset does not contain the birth and death counts, only the population counts at different time points, the distribution needed for a maximum likelihood analysis is that of the difference between two Poisson random variables, known as the Skellam distribution. An analysis comparing the method of moments to maximum likelihood techniques in this setting was done by Alzaid et al. 2010 [32].